Grade 12th passMechanics

Two boys are left in the middle of a frictionless floor. One boy of mass 20 kg gives a blow of 100N, lasting one milli second to the other. He himself will move with a speed of (1) 1m/s(2) 0.5m/s(3) 0.01m/s(4) 0.005m/s

Vikash Kumar Singh
8 Years agoGrade 12th pass

1 Answer

Vishwas prajapati
8 Years ago
Two boys are left in the middle of a fritionlessfloor. Therefore,Mass=20kgForce=100mTime= 10^-3F=mv/tV=ft/m (arrange)V=100×10^-3/20V=0.005m/sHence ,he himself will move with a speed of 0.005 m/s